OpenStudy (anonymous):

Which is a true statement about a 45-45-90 triangle? A. The hypotenuse is times as long as either leg. B. Each leg is times as long as the hypotenuse. C. The hypotenuse is times as long as either leg. D. Each leg is times as long as the hypotenuse.

OpenStudy (anonymous):

need help

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

a^2 + b^2 = c^ a = one side b = other side c = hypotenuse

OpenStudy (anonymous):

c^2

OpenStudy (anonymous):

so if 2^2 + 2^2 = sqrt 8

OpenStudy (anonymous):

okay so C correct

OpenStudy (anonymous):

each side = 2 hypotenuse = 2.82 i dont think any of your options fit, unless it makes sense to you.

OpenStudy (anonymous):

yea the correct anwser was C

OpenStudy (anonymous):

ok, good luck with your study.